I need an equation that models how old in years each of you will be, when your ages add up to 150 years old. The ages are 42, 72, and 15
I cant seem to get make an equation. However, I think the answer is 49, 79, 22... Cand you help me get an equation?

Answer by mananth(16946)

I need an equation that models how old in years each of you will be, when your ages add up to 150 years old. The ages are 42, 72, and 15
Ages = 42, 72,15 now
let the sum of the ages be 150 after x years
..
their ages then will be x+42, x+72, x+15
..
x+42+x+72+x+15=150
3x+129=150
add -129
3x+129-129=150-129
3x=21
divide by 3
x= 7..
Their ages will be 42+7 =49, 72+7=79, 15+7=22 years